# Application of Haldane’s statistical correlation theory in classical systems

### v2

@@ IV Extended non-linear models
V Conclusion and Outlook
In this work, I have given clear-cut evidence that the special constraint used in the purely classical derivation of the CFES statistics is indeed linked to the indistinguishability of the system. Usually, the linear form of Haldane’s theory is used in deriving QFES-type equations. The non-linear modification introduced in this letter helps us to understand the relationship between statistical correlation and indistinguishability using the zero-moment theorem. For a self-correlating system, a quasi-classical derivation using the non-linear statistical correlation theory directly produces CFES-equivalent intermediate statistics. Thus, I have shown that Haldane’s statistical correlation theory lies at the core of CFES statistics.
I have shown that one can derive many intermediate statistical models using appropriate choices for the $b_{l}$ coefficients, which will obey Eq. 17 and produce physically meaningful population distribution. An interesting question can be asked: Is self-correlation necessary to show quantum effects? Currently, I am exploring a binary inter-correlating (not self-correlating) system, which can answer this question and produce interesting results. The application of CFES statistics in enhanced sampling methods in molecular dynamics simulations of classical systems is also being investigated.
Acknowledgement
I am thankful to the National Institute of Technology, Tiruchirappalli, for providing the necessary funding for this research and to Prof. Sanjib Senapati for useful discussions.

### v3

@@ 4 Extended Non-Linear Models
5 Conclusion and Outlook
The inter-dependency of statistical correlation and indistinguishability is an interesting aspect of this letter. Introducing a non-linear modification to the Haldane’s theory of statistical correlation, I have shown that the special constraint used in the purely classical derivation of the CFES statistics is indeed linked to the indistinguishability of the particles. For a self-correlating system, by inserting the modified Haldane’s theory into classical MB statistics directly produces CFES-equivalent intermediate statistics. Thus, I have shown that Haldane’s statistical correlation theory lies at the core of CFES statistics.
I have shown that one can derive many intermediate statistical models using appropriate choices for the $b_{l}$ coefficients, which will obey Eq. 18 and produce physically realistic population distribution. An interesting question can be asked: Is self-correlation necessary to show quantum effects? Currently, I am exploring a binary inter-correlating (not self-correlating) system, which can answer this question and produce interesting results. The application of CFES statistics in enhanced sampling methods in molecular dynamics simulations of classical systems is also being investigated.
Acknowledgements.
I am thankful to the National Institute of Technology, Tiruchirappalli, for providing the necessary funding for this research.